# Distributed Model Checking in Graphs Classes

### v1

@@ 1.2.2 Counting and Optimization
For establishing Theorem 2, we aim at computing the number of $k$-tuples of vertices $(v_{1},\dots,v_{k})$ for which a quantifier-free FO formula $\varphi(x_{1},\dots,x_{k})$ with $k$ free variables holds. let $G=(V,E)$ be the input graph, and $L:V\to 2^{\Lambda}$ be its labeling. We stress the fact that the formula $\varphi$ is assumed to be quantifier-free, and, in particular, it has no other variables than $(x_{1},\dots,x_{k})$. As for model checking, we use the fact that there exists an integer function $g$ depending only on the class $\mathcal{G}$ such that, for any graph $G\in\mathcal{G}$, and any constant $p$, $G$ admits a coloring with $g(p)$ colors such that any subset of at most $p$ colors induces a subgraph of treedepth at most $p$. Thanks to [68], such a coloring can be computed in $\mathcal{O}(\log n)$ rounds in CONGEST. For $p=k$, each $k$-uple of vertices $(v_{1},\dots,v_{k})$ of the labeled graph in $(G,L)$ satisfying the formula $\varphi(x_{1},\dots,x_{k})$ covers some set $S$ of colors, with $|S|\leq p$. Assuming given an $\mathcal{O}(D+\log n)$-round algorithm counting the solutions of $\varphi(x_{1},\dots,x_{k})$ on labeled subgraphs $(G[S],L)$ of treedepth at most $k$, we show how to count the solutions on the whole labeled graph $(G,L)$. For any two sets of colors $S_{1},S_{2}$, the set of solutions of $\varphi(x_{1},\dots,x_{k})$ on $(G[S_{1}\cap S_{2}],L)$ is exactly the intersection of the set of solutions for $(G[S_{1}],L)$ and for $(G[S_{2}],L)$. This holds thanks to the fact that $\varphi(x_{1},\dots,x_{k})$ has no variables other than $(x_{1},\dots,x_{k})$ — this is the reason why our technique cannot be extended to arbitrary formulas. As a consequence, by counting the solutions on $(G[S],L)$ for every set $S$ of at most $k$ colors and then using inclusion-exclusion principle, we are able to counting solutions of $\varphi$ on $(G,L)$.
@@
For establishing Theorem 3, we first use the whole machinery developed for the model-checking (cf. Theorem 1) for transforming the initial formula $\varphi(\overline{x})$ and the labeling $L$ of $G$ into an existential formula $\widehat{\varphi}(\overline{x})=\exists{\overline{y}}\psi(\overline{x},\overline{y})$ and a labeling $\widehat{L}$ of $G$ such that $\psi$ is quantifier-free, and for any $k$-uple of vertices $(v_{1},\dots,v_{k})$, $(G,L)\models\varphi(v_{1},\dots,v_{k})\iff(\widehat{G},\widehat{L})\models\widehat{\varphi}(v_{1},\dots,v_{k})$. This process can be implemented in $\mathcal{O}(D+\log n)$ rounds in CONGEST. We emphasis that the variables of $\widehat{\varphi}$ are precisely the free variables of $\varphi$, but $\widehat{\varphi}$ may have a new set $\overline{y}$ of existentially-quantified variables. Let $p=|\overline{x}|+|\overline{y}|$. We construct now a coloring of $G$ using $f(p)$ colors in $\mathcal{O}(\log n)$ rounds, and we observe that, for any $k$-uple of vertices $(v_{1},\dots,v_{k})$ of $G$, we have $(G,L)\models\varphi(v_{1},\dots,v_{k})$ if and only if there exists a set $S$ of at most $p$ colors such that $(G[S],\widehat{L})\models\widehat{\varphi}(v_{1},\dots,v_{k})$. Indeed, $S$ is obtained by guessing not only the colors of $(v_{1},\dots,v_{k})$, but also the colors of the tuple of vertices assigned to $\overline{y}$. Optimizing the solution of $\varphi$ on $(G,L)$ boils down to optimizing over all optimal solutions of $\widehat{\varphi}$ on graphs $(G[S],\widehat{L})$ of treedepth at most $p$.
@@
Based on the above, it remains to deal with optimization and counting for FO formulas on labeled subgraphs $G[S]$ of $G$, each of constant treedepth. For this purpose we extend the results from [28] stating that counting and optimization of FO formulae can be achieved on connected labeled graphs of bounded treedepth, in constant number of rounds. Indeed, we show that one can perform optimization and counting for FO formulae over an induced — not necessarily connected — subgraph of constant treedepth of the labeled graph $G$, in $\mathcal{O}(D)$ rounds. In [28], the algorithm starts by constructing a tree decomposition of small width of the graph, and then performs dynamic programming over this tree decomposition. If the input graph of small treedepth is connected, this directly enables to construct a rooted tree decomposition of constant depth and constant width. Moreover, the decomposition tree is a spanning tree of the graph, which allows communications to proceed with dynamic programming bottom-up the tree in a constant number of rounds. The root of the tree can then compute the output by simulating the sequential algorithm for optimization and counting on graphs of bounded treewidth of [5]. However, in this paper, the graphs $G[S]$ may not be connected, and thus we have to deal with a forest decomposition $F$ instead of a mere tree decomposition. A breadth-first search starting from an arbitrary root vertex can reach all the roots of the trees in the forest $F$, resulting in a rooted tree of depth $D+\mathcal{O}(1)$. This tree can serve as a base for constructing a tree-decomposition of $G[S]$. Moreover, since the nodes of the tree decomposition correspond to vertices of $G$, the sequential dynamic programming algorithm of [5] can be simulated for counting and optimization on $G[S]$, in $D+\mathcal{O}(1)$ rounds. The information stored in the dynamic programming tables can indeed be computed in a sequence of rounds, starting from the leaves, upward to the root of the tree decomposition by communicating along the edges of $G$.
Altogether, the whole process for counting and optimization for formula $\varphi$ on the pair $(G,L)$ costs $\mathcal{O}(D+\log n)$ rounds in CONGEST.
1.2.3 Distributed Certification
At a high lever, our certification scheme for an FO formula $\varphi$ and a class of graphs $\mathcal{G}$ with bounded expansion essentially consists for the prover to provide each node with the transcript of the execution our distributed model-checking algorithm at this node. The role of the verifier is then to check (in a distributed manner) whether this transcript encodes a correct execution of our algorithm. Note that the certification scheme does not have to certify the bounded expansion property of $\mathcal{G}$, which is a promise. On the other hand, given coloring of the input graph $G=(V,E)\in\mathcal{G}$, the certification scheme must certify that any set $S$ of $k=|\varphi|$ colors does induce a subgraph with treedepth $k$. This is done in an ad hoc manner, and not by providing each node with a transcript of the coloring algorithm from [68]. The main difficulty is to encode the transcript of the quantifier elimination procedure, i.e., the transformation from $(\varphi,G,L)$ to $(\widehat{\varphi},\widehat{L}_{\varphi})$ with $\widehat{\varphi}$ quantifier-free.
[…]
$(\sharp\mbox{red}=\top\land c(r)\geq 1)\lor(\sharp\mbox{red}=\bot\land c(r)<1).$
If any of these test fails at a node, this node rejects, otherwise it accepts. What remains is to certify the spanning tree $T$, which is standard technique using $\mathcal{O}(\log n)$-bit certificates [57].
@@ 1.2.3 Distributed Certification
Note that, at a first glance, it may seem over complicated to use quantifier elimination for certifying a formula of the type $\varphi(\bar{x})=\exists y\,\psi(\bar{x},y)$. Indeed, to certify that $\varphi(\bar{x})=\top$, it may seem sufficient to certify the existence of a vertex $y$ that yield $\psi(\bar{x},y)=\top$, which can be done easily thanks to a spanning tree rooted at $y$, consuming $\mathcal{O}(\log n)$-bit certificates [57]. The situation is however more complex than it seems, for at least two reasons. First, in the intermediate steps $\psi$ may refers to free variables $(x_{1},\dots,x_{\ell})=\bar{x}$, which makes the identification of a vertex $y$ such that $\psi(\bar{x},y)=\top$ non-trivial. Second, there may be negations in the global formula, that is, one may have to consider also the case where $\varphi(\bar{x})=\neg\exists y\,\psi(\bar{x},y)$, and certifying the non existence of a vertex $y$ is not straightforward. A typical example is the case of certifying the existence of a 4-cycle, which is trivial using $\mathcal{O}(\log n)$-bit certificate, in contrast to certifying the non existence of a 4-cycle, which requires certificates on $\tilde{\Omega}(\sqrt{n})$ bits in general graphs.
@@ 1.3.1 Distributed Computing on Sparse Networks
The quest for efficient (sublinear) algorithms for solving classical graph problems in the CONGEST model dates back to the seminal paper by [37], which presented an algorithm for constructing a Minimum Spanning Tree (MST). Since then, a long series of problems have been addressed, including connectivity decomposition by [7], tree embeddings by [44], $k$-dominating set by [60], Steiner trees by [61], min-cut by [43, 63], max-flow by [42], and shortest paths by [53, 62], among others. Additionally, algorithms tailored to specific classes of networks have been developed, such as depth-first search (DFS) for planar graphs [45], MST for bounded genus graphs [50], and maximal independent set (MIS) for networks excluding a fixed minor [8].
Overcoming the $\Omega(D+\sqrt{n})$ barrier has been a significant focus in this field. Notably, the work by [39, 40] showed a planar embedding in $\mathcal{O}(D\cdot\min{\log n,D})$, while [50] demonstrated that an embedding is unnecessary for MST. Later, these results were extended to graphs of bounded genus by [51], and graphs excluding a fixed graph as a minor by [52, 41]. Furthermore, algorithms achieving $\tilde{\mathcal{O}}(D)$ complexity in the CONGEST model for bounded treewidth graphs have been developed for various problems, including matching, the directed single-source shortest paths problem, bipartite unweighted maximum matching, and girth [54].
@@ 1.3.2 Centralized Meta-Theorems
From the technical perspective, the most important for our work is the quantifier elimination procedure introduced in [16] and in [48], see also [71]. Similar to these works, we apply the idea of modifying the structure by adding new unary relations and functions without changing the Gaifman graph.
1.3.3 Distributed Meta-Theorems
We are not aware of any meta-theorem combining logical and combinatorial constraints within the CONGEST model, with one exception. The very recent paper by [28] establishes that every MSO formula on graphs with bounded treedepth is decidable in a constant number of rounds within the CONGEST model. The results of [28] are incomparable withTheorem 1. Indeed, Theorem 1 applies to graph classes that are significantly broader than graphs of bounded treedepth, but the number of rounds of the algorithm by [28] is independent of the number of vertices of the input graph. We incorporate some of the methods developed by [28] in our paper.
Another general result for CONGEST model is the work of [68], who gave a distributed algorithm for computing a low treedepth decomposition of graphs of bounded expansion, running in $O(\log n)$ rounds. We use this construction in our algorithm as a subroutine. In the same paper, Nešetřil and Ossona de Mendez observe that their algorithm could be used to check whether a graph of bounded expansion contains a given connected subgraph $H$ of constant size, in $O(\log n)$ rounds.
Distributed certification is a very active area of research, with numerous results emerging since the survey by [23]. Recent work has focused on approximate variants of the problem, similar to property testing [19, 24], and [17]. Notably, it has been shown that every monotone (i.e., closed under taking subgraphs) and summable (i.e., stable under disjoint union) property $\Pi$ has a compact approximate certification scheme in any proper minor-closed graph class [21]. Other recent contributions have explored augmenting the power of the verifier in certification schemes. This includes tradeoffs between the size of the certificates and the number of rounds of the verification protocol by [26], randomized verifiers by [33], quantum verifiers by [29], and several hierarchies of certification mechanisms. These mechanisms include games between a prover and a disprover [1, 25], interactive protocols [11, 55, 64], zero-knowledge distributed certification [3], and distributed quantum interactive proofs [36].
Several meta-theorems are known for distributed certification. A significant result in the field was achieved in [22], who demonstrated that any MSO formula can be locally certified on graphs with bounded treedepth using a logarithmic number of bits per node, representing a gold standard in certification. This theorem has numerous implications for certification; for more details, we refer to [22]. Notably, the FO property of $C_{4}$-freeness and the MSO property of non-3-colorability, which typically require polynomial-size certificates in general graphs, can be certified with just $\mathcal{O}(\log n)$-bit certificates in graphs of bounded treedepth. The result by Bousquet et al. has been extended to more comprehensive classes of graphs, including graphs excluding a small minor [6], graphs of bounded treewidth, and graphs of bounded cliquewidth. However, this extension comes at the cost of slightly larger certificates of $\mathcal{O}(\log^{2}n)$ bits, as shown in [32] and [31], respectively.

### v2

@@ 1.3.4 Model Checking, Counting, Optimization, and Certification
  • Quantifier-free formulas: As for model checking, we use the fact that there exists an integer function $g$ depending only on the class $\mathcal{G}$ such that, for any graph $G\in\mathcal{G}$, and any constant $p$, $G$ admits a coloring with $g(p)$ colors such that any subset of at most $p$ colors induces a subgraph of treedepth at most $p$. For $p=k$, each $k$-uple of vertices $(v_{1},\dots,v_{k})$ of the labeled graph in $(G,\ell)$ satisfying the formula $\varphi(x_{1},\dots,x_{k})$ covers some set $U$ of colors, with $|U|\leq p$. Assuming given an $\mathcal{O}(D+\log n)$-round algorithm counting the solutions of $\varphi(x_{1},\dots,x_{k})$ on labeled subgraphs $(G[U],\ell)$ of treedepth at most $k$, we show how to count the solutions on the whole labeled graph $(G,\ell)$. For any two sets of colors $U_{1},U_{2}$, the set of solutions of $\varphi(x_{1},\dots,x_{k})$ on $(G[U_{1}\cap U_{2}],\ell)$ is exactly the intersection of the set of solutions for $(G[U_{1}],\ell)$ and for $(G[U_{2}],L)$. This holds thanks to the fact that $\varphi(x_{1},\dots,x_{k})$ has no variables other than $(x_{1},\dots,x_{k})$. As a consequence, by counting the solutions on $(G[U],\ell)$ for every set $U$ of at most $k$ colors, and then using the inclusion-exclusion principle, we are able to count solutions of $\varphi$ on $(G,\ell)$.
  • General formulas: In order to use the previous stage on quantifier free formulas, we first proceed with quantifier elimination. This introduces new variables and new predicates, especially lca predicates, which prevents us from directly applying the previous stage. To overcome this difficulty, we eliminate the lca predicates by introducing new variables, which are existentially quantified. To avoid over-counting, we enforce unicity, i.e., at most one instance of each newly introduced variable may satisfy the formula. Informally, it is as if every existential quantifier “there exists $y$ such that” was replaced by “there exists a unique $y$ such that”.
@@
To establish Theorem 1.4, we first utilize the entire quantifier elimination machinery developed for distributed model-checking in CONGEST. We emphasize that, during the process, the variables of the quantifier-free formula $\widehat{\varphi}$ are precisely the free variables of $\varphi(\bar{x})$, but $\widehat{\varphi}$ may have a new set $\overline{y}$ of existentially-quantified variables. Let $p=|\overline{x}|+|\overline{y}|$. We construct now a coloring of $G$ using $g(p)$ colors in $\mathcal{O}(\log n)$ rounds, and we observe that, for any $k$-uple of vertices $(v_{1},\dots,v_{k})$ of $G$, we have $(G,\ell)\models\varphi(v_{1},\dots,v_{k})$ if and only if there exists a set $U$ of at most $p$ colors such that $(G[U],\widehat{\ell})\models\widehat{\varphi}(v_{1},\dots,v_{k})$. Indeed, $U$ is obtained by guessing not only the colors of $v_{1},\dots,v_{k}$, but also the colors of the tuple of vertices $u_{1},\dots,u_{p-k}$ assigned to $\overline{y}$. Optimizing the solution of $\varphi$ on $(G,\ell)$ boils down to optimizing over all optimal solutions of $\widehat{\varphi}$ on graphs $(G[U],\widehat{\ell})$ of treedepth at most $p$.
@@
Based on the above, it remains to deal with optimization and global counting for FO formulas on labeled subgraphs $G[U]$ of $G$, each of which has a constant treedepth. For this purpose, we extend the results from [25] stating that counting and optimization of FO formulae can be achieved on connected labeled graphs of bounded treedepth, in a constant number of rounds. Indeed, we show that one can perform optimization and counting for FO formulae over an induced (not necessarily connected) subgraph of constant treedepth of the labeled graph $G$, in $\mathcal{O}(D)$ rounds. In [25], the algorithm starts by constructing a tree decomposition of small width of the graph, and then performs dynamic programming over this tree decomposition. If the input graph of small treedepth is connected, this directly enables the construction of a rooted tree decomposition of constant depth and constant width. Moreover, the decomposition tree is a spanning tree of the graph, which allows communications to proceed with dynamic programming bottom-up the tree in a constant number of rounds. The root of the tree can then compute the output by simulating the centralized algorithm for optimization and counting on graphs of bounded treewidth of [6]. However, in our case, the graphs $G[U]$ may not be connected, and thus we must deal with a forest decomposition $F$ instead of a mere tree decomposition. A BFS tree rooted at an arbitrary vertex can span all the roots of the trees in the forest $F$, resulting in a rooted tree of depth $D+\mathcal{O}(1)$. This tree can serve as a base for constructing a tree-decomposition of $G[U]$. Moreover, since the nodes of the tree-decomposition correspond to vertices of $G$, the centralized dynamic programming algorithm of [6] can be simulated for counting and optimization on $G[U]$, in $D+\mathcal{O}(1)$ rounds. The information stored in the dynamic programming tables can indeed be computed in a sequence of rounds, starting from the leaves, upward to the root of the tree decomposition by communicating along the edges of $G$. Altogether, the whole process for counting and optimization for formula $\varphi$ on the pair $(G,\ell)$ costs $\mathcal{O}(D+\log n)$ rounds in CONGEST.
Our method for local counting follows the same lines as that for global counting. That is, we first reduce the study to quantifier-free formulas, using the (locality preserving) quantifier elimination method. Similarly to the case of global counting, this reduction introduces lca-predicates, which need to be eliminated in a locality-preserving fashion. It turns out that the method applied to the case of global counting happens to preserve locality, although the proof is somewhat technical. For quantifier-free formulas, we essentially apply the same idea as for global counting, based on the inclusion-exclusion principle. However, we also use the fact that the formula is local, by considering some set of colors $U$, and a tuple of vertices $\overline{v}$ such that $(G[U],\ell)$ satisfies $\varphi(\overline{v})$. Then, since $\varphi$ is local, all vertices in $\overline{v}$ belong to the same connected component of $G[U]$, which implies that we only need to count the solutions to $\varphi$ in the connected components of $G[U]$. The latter are connected graphs of bounded treedepth, and thus counting is achieved in $\mathcal{O}(1)$ rounds using the algorithm in [25].
Distributed Certification.
@@
Let us start with a preliminary remark. At first glance, it may seem overcomplicated to use quantifier elimination for certifying a formula of the type $\varphi(\bar{x})=\exists y\,\psi(\bar{x},y)$. Indeed, to certify that $\varphi(\bar{x})=\top$, it may seem sufficient to certify the existence of a vertex $y$ that yield $\psi(\bar{x},y)=\top$, which can be done easily thanks to a spanning tree rooted at $y$, consuming $\mathcal{O}(\log n)$-bit certificates [48]. The situation is, however, more complex than it seems, for at least two reasons. First, in the intermediate steps $\psi$ may refers to free variables $(x_{1},\dots,x_{\ell})=\bar{x}$, which makes the identification of a vertex $y$ such that $\psi(\bar{x},y)=\top$ non-trivial. Second, there may be negations in the global formula, that is, one may have to consider also the case where $\varphi(\bar{x})=\neg\exists y\,\psi(\bar{x},y)$, and certifying the non existence of a vertex $y$ is not straightforward. A typical example is the case of certifying the existence of a 4-cycle, which is trivial using $\mathcal{O}(\log n)$-bit certificate, in contrast to certifying the non existence of a 4-cycle, which requires certificates on $\tilde{\Omega}(\sqrt{n})$ bits in general graphs.
At a high level, our certification scheme for an FO formula $\varphi$ and a class of graphs $\mathcal{G}$ with bounded expansion essentially consists of the prover providing each node with the transcript of the execution of our distributed model-checking algorithm at this node. The role of the verifier is then to check (in a distributed manner) whether this transcript encodes a correct execution of our algorithm. Note that the certification scheme does not have to certify the bounded expansion property of $\mathcal{G}$, which is a promise. However, given a coloring of the input graph $G=(V,E)\in\mathcal{G}$, the certification scheme must certify that any set $U$ of $k=|\varphi|$ colors does induce a subgraph with treedepth $k$. This is done in a specific manner, using an algorithm by Bousquet et al. [7] for certifying MSO formulas on graphs of bounded treedepth with certificates of size $\mathcal{O}(\log n)$. The main difficulty is to encode the transcript of the quantifier elimination procedure, i.e., the transformation from $(\varphi,G,\ell)$ to $(\widehat{\varphi},\widehat{\ell})$ with $\widehat{\varphi}$ quantifier-free. Each lca predicate refers to a local condition, and the value of such a predicate can be certified locally at each node, in forests of bounded depth. On the other hand, the unary predicates may refer to non-local properties. Nevertheless, we demonstrate that verifying the correctness of the unary predicates can be accomplished using a spanning tree whose root is capable of making decisions. (Certify a spanning tree is standard using $\mathcal{O}(\log n)$-bit certificates [48].)
@@ 1.4 Related Work
The results of our paper lie at the intersection of two research directions. The first is the study of distributed computing and distributed certification schemes on sparse networks, such as planar graphs, or graphs excluding a fixed minor, for fundamental graph optimization problems. The second is the development of meta-theorems for sparse graph classes in centralized computation.
@@ 1.4.1 Distributed Computing on Sparse Networks
One of the popular areas of research in distributed computing concerns the study of classical combinatorial problems on sparse networks, including planar graphs, bounded-degree graphs, graphs of bounded treewidth, and graphs excluding a fixed minor. In such settings, a central goal is to develop distributed algorithms that exploit structural sparsity to achieve improved efficiency—compared to general networks—particularly in terms of communication rounds. Typical problems studied in this context include coloring, matching, dominating sets, cuts, spanning trees, and network decompositions. Research directions focus on understanding how sparsity affects the complexity of these problems in standard distributed models such as LOCAL and CONGEST, and on designing algorithms with optimal or near-optimal round complexity. This line of work often builds on structural properties of sparse networks, revealing deep and fruitful connections between structural graph theory and distributed computation.
Examples of the work in this direction include distributed algorithms for maximal matchings, maximal independent sets (MIS), and colourings in graphs of bounded vertex degrees [56], depth-first search (DFS) for planar graphs [34], minimum spanning tree (MST) and min-cut for planar, bounded genus, and $H$-minor-free graphs [31, 32, 40, 41, 42, 33], minimum dominating set on planar and $H$-minor-free graphs [1, 4, 12, 13, 52], and MIS for graphs excluding a fixed minor [8].
The round complexity $\tilde{\mathcal{O}}(D)$ achieved in Theorem 1.2, is considered a “gold standard” for algorithms on sparse graphs in the CONGEST model. It has been obtained, for instance, for the Minimum Spanning Tree (MST) problem on planar graphs and, more generally, on $H$-minor-free graphs [31, 32, 40, 41, 42, 33]. Algorithms with $\tilde{\mathcal{O}}(D)$ complexity in the CONGEST model have also been developed for bounded-treewidth graphs for a variety of problems, including matching, the directed single-source shortest paths problem, bipartite unweighted maximum matching, and girth [46].
We are not aware of any meta-theorem that combines logical and combinatorial constraints within the CONGEST model, with one exception. The paper [25] establishes that every MSO formula on graphs of bounded treedepth is decidable in a constant number of rounds in the CONGEST model. The results of [25] are incomparable with Theorems 1.1 and 1.2. Indeed, Theorems 1.1 and 1.2 apply to graph classes that are significantly broader than graphs of bounded treedepth, whereas the number of rounds in the algorithm of [25] is independent of the number of vertices in the input graph. We incorporate some of the methods developed in [25] in our work.
Regarding distributed algorithms on graphs of bounded expansion, the foundational work in this area is due to Nešetřil and Ossona de Mendez [55]. In the CONGEST model, they presented a distributed algorithm for computing a low treedepth decomposition of graphs of bounded expansion in $O(\log n)$ rounds. We use this construction as a subroutine in our algorithms. In the same paper, Nešetřil and Ossona de Mendez also used this decomposition to design an algorithm for checking whether a graph of bounded expansion contains a given connected subgraph $H$ of constant size, also in $O(\log n)$ rounds. Theorem 1.1 significantly extends this result.
Approximation algorithms for minimum dominating set on graphs of bounded expansion in LOCAL model were developed in [1, 44, 50]. In [62], Siebertz and Vigny discuss the complexity of the model-checking problem for first-order logic on graphs of bounded expansion, in the CONGESTED CLIQUE model.
Distributed certification is a very active area of research, see [23] for a survey. Several meta-theorems are known for distributed certification. A significant result in the field was achieved in [22], who demonstrated that any MSO formula can be locally certified on graphs with bounded treedepth using a logarithmic number of bits per node, representing a gold standard in certification. This theorem has numerous implications for certification; for more details, we refer to [22]. Notably, the FO property of $C_{4}$-freeness and the MSO property of non-3-colorability, which typically require polynomial-size certificates in general graphs, can be certified with just $\mathcal{O}(\log n)$-bit certificates in graphs of bounded treedepth. The result by Bousquet et al. has been extended to more comprehensive classes of graphs, including graphs excluding a small minor [7], graphs of bounded treewidth [2, 10, 27], and graphs of bounded cliquewidth [26].
@@ 1.4.2 Centralized Meta-Theorems
From a technical perspective, the most important aspect for our work is the quantifier elimination procedure introduced in [18] and [38], as well as in [59]. Similar to these works, we apply the idea of modifying the structure by adding new unary relations and functions without changing the Gaifman graph.